OpenStudy (bekkah323):

2.16 is 4% of what number?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

4% = 4/100 2.16 = (4/100) x solve for x :)

OpenStudy (bekkah323):

54?

OpenStudy (skullpatrol):

Correct :)